# How Worst-Case Are Adversarial Attacks? Linking Adversarial and Statistical Robustness

### v1

@@ 6 Experiments
In this section, we benchmark adversarial attacks using the proposed analysis across different architectures and settings, and study attack effectiveness, statistical behavior, and model confidence under noisy risk evaluation. Finally, we present ablation studies of the proposed attack.
@@ 6.1 Experimental setup
Attacks selected. In addition to the proposed DN attack, we considered a subset of common and widely used adversarial methods that well categorize distinct optimization behaviors. PGD (Madry et al., 2018) performs a direct steepest-ascent optimization to maximize the loss within the admissible perturbation set. While highly effective in worst-case settings, PGD and related strong attacks (e.g., CW (Carlini & Wagner, 2017), DF (Moosavi-Dezfooli et al., 2016)) primarily emphasize decision-boundary crossing rather than neighborhood exploration (less aligned with the objectives of this noisy risk analysis). To explicitly account for stochastic optimization effects, we include MI-FGSM (Dong et al., 2018) and PGN (Ge et al., 2023). MI-FGSM introduces momentum as a form of structured noise. PGN further extends this by injecting noise at each iteration and incorporating Hessian-based regularization. Finally, we consider the simple FGSM (Szegedy et al., 2014), which, despite its simplicity, exhibits informative behavior under this analysis.
Regarding attack settings, we consider both $\ell_{\infty}$ and $\ell_{2}$ threat models. For attacks that use momentum, we adopt standard parameters with momentum $\mu=1$, step size $\alpha=\varepsilon/N$, number of steps equal to $10$, and $N=10$ for both DN and PGN. For PGD, we use a step size $\alpha=2/255$ for $\ell_{\infty}$ attacks and $\alpha=0.25$ for $\ell_{2}$ attacks. PGD is evaluated with both 5 and 20 steps. Finally, for DN we used $\kappa_{\text{adv}}=100$, and conduct ablation studies in Section 6.5.
Metrics. The directional noisy risk error $\mathcal{R}$ (Eq. 4) was estimated via a Monte Carlo approach using $n=256$ samples. Since $\mathcal{R}$ is defined at the sample level, we also report its average over the test set, denoted by $\bar{\mathcal{R}}$. Additionally, we report two complementary indicators: $\kappa_{\bar{\mathcal{R}}=0.25}$ and $\bar{\mathcal{R}}_{\kappa^{*}}$, where $\kappa^{*}=D^{1/4}$. The former can be interpreted as a plausibility risk score, defined as the smallest concentration parameter $\kappa$ such that $\bar{\mathcal{R}}\geq 0.25$, with lower values indicating failure regions closer to noise-like perturbation distributions. The latter, $\bar{\mathcal{R}}_{\kappa^{*}}$, measures the noisy risk at the reference concentration $\kappa^{*}$ (see Sec. 4 and Sec. 6.3), corresponding to success rate that remain statistically plausible under uniform noise on the $\ell_{p}$-sphere.
@@
Figure 3: The average risk-aware error $\bar{\mathcal{R}}$ (y-axis) is shown as a function of $\kappa$ for different attacks under $\ell_{\infty}$ (first and second rows, $\varepsilon=8/255$ and $4/255$) and $\ell_{2}$ (third and fourth, $\varepsilon=0.5$ and $\varepsilon=1.0$) threat models, evaluated on ResNet-50, ViT-B, WideResNet-101, and VGG-16. The reference value $\kappa^{*}$ is indicated by dotted lines, highlighting a statistically plausible risk regime.
@@ 6.2 Benchmarking Adversarial Attacks
To enable a meaningful comparison of adversarial perturbations under the noisy risk perspective, Table 1 reports a benchmark of selected attacks based on the proposed metrics, while also including the classical attack success rate (ASR in table) for comparison. Results are reported for both $p=\infty$ and $p=2$ across different $\varepsilon$. Best and second-best are highlighted using boldface and underlining, respectively.
In all configurations, the proposed DN attack consistently achieves the highest directional noisy risk error in low regimes $\kappa$, as indicated by $\bar{\mathcal{R}}_{\kappa^{*}}$. FGSM often emerges as the second-best method at $\kappa^{*}=D^{1/4}$, despite its simplicity and the fact that it is commonly regarded in the literature as a weak attack. Similar consideration also could be reported for the $\kappa_{\bar{\mathcal{R}}=0.25}$ indicator, showing the benefits of the above mentioned approaches for low $\kappa$ values. Conversely, this comparison highlights a key limitation of stronger techniques: attacks such as PGD, while highly effective in the pointwise limit (as shown by the ASR, i.e., $\bar{\mathcal{R}}$ for $\kappa\rightarrow\infty$), do not necessarily correspond to representative or stable failure modes when noisy perturbations are taken into account. Consequently, ASR alone can be misleading, and even simple attacks such as FGSM may be more informative than stronger iterative ones when evaluated through a noisy risk lens, as shown by the better results achieved in both the two metrics proposed. Interestingly, PGN exhibits limited noisy risk scores: despite outperforming PGD in the proposed analysis, its Hessian regularization, known to enhance transferability across achitectures (Ge et al., 2023), is effective mainly in high-$\kappa$ regimes. This questions the connection between stable loss curvature and robustness to statistically broad noisy perturbations.
Figure 5: Correlation between uniform project noise ($\mathcal{R}$ at $\kappa=0$) and $\mathcal{R}$ varying $\kappa$, using DN. (a) per-sample relation on ImageNet-Val samples for $\ell_{\infty}$ with $\varepsilon=16/255$. (b) Spearman correlation $\rho$ vs. $\kappa$ for ImageNet-Val (top: $\varepsilon=16/255$ for $\ell_{\infty}$, $\varepsilon=2.0$ for $\ell_{2}$) and CIFAR-10 (bottom: $\varepsilon=8/255$ for $\ell_{\infty}$, $\varepsilon=1.0$ for $\ell_{2}$)
@@
Importantly, the reference concentration $\kappa^{*}=D^{1/4}$ (vertical dotted lines) reveals a substantial gap between $\bar{\mathcal{R}}$ and the classical attack success rate (ASR). From a probabilistic perspective, this gap highlights that standard adversarial attacks primarily exploit brittle failure modes, which have limited relevance under realistic perturbation distributions (Fawzi et al., 2018; Carlini et al., 2019). Despite this common behavior, the DN attack exhibits a faster increase of $\bar{\mathcal{R}}$ with inital values of $\kappa$, indicating broader, noise-aligned failure regions at moderate concentrations. While not optimized for worst-case ASR, DN and FGSM remain more informative under noisy risk evaluation, with similar, though more limited, trends observed for MIFGSM and PGN due to the use of noise during optimization.
@@ 6.3 Empirical validation of $\kappa^{*}$
The reference concentration $\kappa^{*}=D^{1/4}$ is introduced based on the analysis presented in Section 4 and Appendix B. To further validate this choice from an empirical perspective, we analyze the behavior of the directional noisy risk error for small values of $\kappa$, focusing on DN-based attacks, which, according to the previous benchmarks, achive best performance in low $\kappa$ regimes. Specifically, we study the sample-wise correlation between $\mathcal{R}$ computed at fixed $\kappa$ values and that obtained under uniform projected noise ($\mathcal{R}$ with $\kappa=0$). As illustrated in Figure 5(a), for small concentration values (e.g., $\kappa=2$), the directional noisy risk error remains strongly correlated with the uniform-noise case; however, as $\kappa$ increases (e.g., $\kappa=10$), this similarity progressively decreases, indicating that the attack achieves high success rates only in highly directional regimes, while remaining largely ineffective under uniform noise.
To quantify this trend across multiple samples, Figure 5(b) reports the Spearman rank correlation coefficient $\rho_{s}$ between the $\mathcal{R}$ values at a given $\kappa$ (x-axis) and those at $\kappa=0$. Results are shown for both ImageNet-Val and CIFAR-10, and for $p=\infty$ and $p=2$, highlighting the effect of different input dimensionalities. In all settings, the rank correlation decreases sharply with $\kappa$, confirming that higher concentrations correspond to increasingly localized and less representative perturbation regimes. Importantly, the reference value $\kappa^{*}=D^{1/4}$ ($\approx 19$ for ImageNet, $\approx 7$ for CIFAR-10) lies in a transitional region where rank correlation with uniform noise remains relatively high, yet begins to decay rapidly. This regime represents a meaningful trade-off: it preserves a strong connection to noise-aligned perturbations while avoiding high-$\kappa$ regime where the analysis becomes dominated by corner-case and too worst-case behavior.
@@ 6.4 On the reliability of model confidence
In addition to the analysis presented above, we investigate how model inference behavior relates to robustness across different $\kappa$ regimes, with the goal of identifying common characteristics of samples that are vulnerable to stochastic noise. Figure 7 reports the value of $\mathcal{R}$ for individual samples as a function of the model confidence on the corresponding clean input (x-axis), i.e., $\exp\big(f_{\hat{y}(x)}(x)\big)/{\sum_{c}\exp\big(f_{c}(x)\big)}.$ The analysis is conducted for different concentration values, $\kappa\in\{5,20,50\}$, using the DN attack, illustrating the transition from more noise-representative to highly directional attack regimes. Results for additional models, exhibiting a similar trend, are reported in Appendix D.
For both ResNet50 and ViT-B, a consistent pattern emerges. In the lowest concentration regime ($\kappa=5$), samples exhibiting high noisy risk ($\mathcal{R}\approx 1$) are predominantly associated with low confidence scores. Conversely, samples predicted with high confidence remain largely robust, with $\mathcal{R}$ close to zero, despite being potentially vulnerable under classical worst-case adversarial evaluations. As $\kappa$ increases to moderate values (e.g., $\kappa=20\approx\kappa^{*}$), high noisy risk gradually extends toward samples with higher confidence. This effect is more pronounced for ResNet models, while ViT architectures exhibit more stable reliability, with high-risk samples remaining largely concentrated among low-confidence predictions. Only at relatively large values of $\kappa$ (e.g., $\kappa\geq 50$, where noisy risk becomes less informative) does the distribution of highly attacked samples ($\mathcal{R}$ high) significantly include high-confidence predictions.
The trend suggests that, under statistically plausible perturbation regimes, adversarial vulnerability could be strongly coupled with model uncertainty. This provides a more nuanced perspective on the relationship between confidence and robustness: while confidence is known to be unreliable under worst-case threat models (Grosse et al., 2018; Galil & El-Yaniv, 2021), it could remain informative for identifying samples that are realistically at risk under noise-aligned adversarial perturbations. This observation highlights a potential connection between robustness evaluation and uncertainty estimation, suggesting that uncertainty-aware analyses (Gawlikowski et al., 2023; Oberdiek et al., 2018) should not be dismissed based on their behavior under worst-case adversarial settings, and still play a meaningful role in safety-oriented model assessment.
@@ 7 Conclusion and Discussion
This work examines the role of adversarial attacks in robustness and safety evaluation. We introduce a probabilistic framework that bridges adversarial and noisy robustness through directional noisy risk analysis, utilizing a concentration parameter to characterize the relationship between adversarial directions and noise-aligned perturbations. We then presented a directional noisy attack that explicitly targets statistically representative failure regions. Results show that many standard adversarial attacks achieve high success only in highly worst-case regimes, while simpler or noise-aware strategies can be more informative when robustness is evaluated from a noisy-risk perspective. These findings, rather than discouraging the use of attacks for this purpose, highlight important limitations of relying solely on attack success rates and invite practitioners to complement attack evaluation with noisy-risk metrics.
@@ Appendix A Sampling Directions on $\ell_{p}$-Spheres
Let $\xi\sim\mathcal{N}(0,I_{d})$ be a standard Gaussian vector in $\mathbb{R}^{d}$. We consider perturbations obtained by projection onto the $\ell_{p}$-sphere of radius $r$, $\eta=\Pi_{\|\cdot\|_{p}=r}(\xi):=r\frac{\xi}{\|\xi\|_{p}}.$
@@
Despite this lack of exact uniformity, Gaussian projection remains a practical and effective approximation in high dimensions. Indeed, by concentration of measure, $\|\xi\|_{p}^{p}=\sum_{i=1}^{d}|\xi_{i}|^{p}$ concentrates sharply around its expectation as $d$ increases. Consequently, normalization by $\|\xi\|_{p}$ primarily rescales $\xi$ by an almost deterministic factor, while preserving a diverse set of directions. As a result, $\xi/\|\xi\|_{p}$ provides a stable set of noise-aligned perturbations that is sufficient for estimating noisy-risk probabilities, while retaining exact uniformity in the $\ell_{2}$ case.
Generality of the proposed analysis.
@@ Appendix B On the Scaling of $\kappa$ and $\kappa^{*}$
This places $\kappa^{\star}$ well within the noise-dominated regime, ensuring that samples remain broadly stochastic while introducing a controlled and measurable directional preference. In typical high-dimensional vision settings ($D\approx 10^{4}$$10^{6}$), this choice corresponds to moderate values of $\kappa$ that are empirically well below $\sqrt{D}$, yet sufficient to induce meaningful directional bias. Please note that $\kappa^{\star}=D^{1/4}$ is not intended to be an optimal or universal choice. Rather, it serves as a convenient reference point for the proposed analysis, lying well within the noise-dominated regime while still inducing a measurable directional bias. The conclusions of this work do not rely on the exact value of $\kappa^{\star}$, but instead on the qualitative behavior of the noisy-risk metric across low-$\kappa$ regimes. Similar trends are observed for a broad range of $\kappa$ values satisfying $\kappa\ll\sqrt{D}$, as further supported by the experimental analysis.
Table 2: Robustness scores under $\ell_{\infty}$ attacks ($\varepsilon\in\{4/255,8/255\}$) and $\ell_{2}$ attacks ($\varepsilon\in\{0.5,1.0\}$). Here, $\bar{\mathcal{R}}_{\kappa^{*}}$ denotes the average risk-aware adversarial error on the test set evaluated at $\kappa^{*}=D^{1/4}$, while $\kappa_{0.25}$ denotes the smallest concentration parameter $\kappa$ for which $\bar{\mathcal{R}}\geq 0.25$. ASR denotes the standard attack success rate, reported for comparison.
@@
Figure 9: Comparison of different adversarial attacks on the CIFAR10 tests set measured by the noisy-risk metric $\bar{\mathcal{R}}$ at small values of $\kappa$ for ResNet-32 and VGG-16 under $\ell_{2}$ perturbations. The dotted lines denote the reference point $\kappa^{*}=D^{1/4}$.

### v2

@@ 6 Experiments
In this section, we benchmark adversarial attacks using the proposed analysis across different architectures and settings, and study attack effectiveness, statistical behavior, and model confidence under perturbation risk evaluation. Finally, we present ablation studies of the proposed attack.
@@ 6.1 Experimental setup
Attacks selected. In addition to the proposed DN attack, we consider a subset of common and widely used adversarial methods that well categorize distinct optimization behaviors. PGD (Madry et al., 2018) performs a direct steepest-ascent optimization to maximize the loss within the admissible perturbation set. While highly effective in worst-case settings, PGD and related strong attacks (e.g., CW (Carlini & Wagner, 2017), DF (Moosavi-Dezfooli et al., 2016)) primarily emphasize decision-boundary crossing rather than neighborhood exploration, less aligned with the objectives of this paper. To explicitly account for stochastic optimization effects, we include MI-FGSM (Dong et al., 2018) and PGN (Ge et al., 2023). MI-FGSM introduces momentum as a form of structured noise. PGN further extends this by injecting noise at each iteration and incorporating Hessian-based regularization. Finally, we consider the simple FGSM (Szegedy et al., 2014), which, despite its simplicity, exhibits informative behavior under our analysis.
Regarding attack settings, we consider both $\ell_{\infty}$ and $\ell_{2}$ threat models. For attacks that use momentum, we adopt standard parameters with momentum $\mu=1$, step size $\alpha=\varepsilon/T$, number of steps $T=10$, and $N=10$ for both DN and PGN. For PGD, we use a step size $\alpha=2/255$ for $\ell_{\infty}$ attacks and $\alpha=0.25$ for $\ell_{2}$ attacks, and we evaluate it with both $5$ and 20 steps. Finally, for DN we used $\kappa_{\text{adv}}=100$, and conduct ablation studies in Section 6.5.
Metrics. The directional perturbation risk $\mathcal{R}$ (Eq. 4) is estimated via a Monte Carlo approach using $n=256$ samples. Since $\mathcal{R}$ is defined at the sample level, we also report its average over the test set, denoted by $\bar{\mathcal{R}}$. Additionally, we report two complementary indicators: $\kappa_{\bar{\mathcal{R}}=0.25}$ and $\bar{\mathcal{R}}_{\kappa^{*}}$, where $\kappa^{*}=D^{1/4}$. The former can be interpreted as a plausibility risk score, defined as the smallest concentration parameter $\kappa$ such that $\bar{\mathcal{R}}\geq 0.25$, with lower values indicating that the attack recognizes more statistically relevant failure regions. The latter, $\bar{\mathcal{R}}_{\kappa^{*}}$, measures the perturbation risk at the reference concentration $\kappa^{*}$ (see Sec. 4 and Sec. 6.3), corresponding to success rate that remain statistically plausible under uniform noise on the $\ell_{p}$-sphere.
@@
Figure 3: The average perturbation risk $\bar{\mathcal{R}}$ (y-axis) is shown as a function of $\kappa$ for different attacks under $\ell_{\infty}$ (first and second rows, $\varepsilon=8/255$ and $4/255$) and $\ell_{2}$ (third and fourth, $\varepsilon=0.5$ and $\varepsilon=1.0$) threat models, evaluated on ResNet-50, ViT-B, WideResNet-101, and VGG-16. The reference value $\kappa^{*}$ is indicated by dotted lines, highlighting a statistically plausible risk regime.
@@ 6.2 Benchmarking Adversarial Attacks
To enable a meaningful comparison of adversarial attacks under the perturbation risk perspective, Table 1 reports a benchmark of selected attacks based on the proposed metrics, while also including the classical attack success rate (ASR in table) for comparison. Results are reported for both $p=\infty$ and $p=2$ across different $\varepsilon$. Best and second-best are highlighted using boldface and underlining, respectively.
In all configurations, the proposed DN attack consistently achieves the highest directional perturbation risk in low regimes $\kappa$, as indicated by $\bar{\mathcal{R}}_{\kappa^{*}}$. FGSM often emerges as the second-best method at $\kappa^{*}=D^{1/4}$, despite its simplicity and the fact that it is commonly regarded in the literature as a weak attack. Similar consideration also could be reported for the $\kappa_{\bar{\mathcal{R}}=0.25}$ indicator, showing the benefits of the above mentioned approaches for low $\kappa$ values. Conversely, this comparison highlights a key limitation of stronger techniques: attacks such as PGD, while highly effective in the pointwise limit (as shown by the ASR, i.e., $\bar{\mathcal{R}}$ for $\kappa\rightarrow\infty$), do not necessarily correspond to representative or stable failure modes when stochastic perturbations are taken into account. Consequently, ASR alone can be misleading, and even simple attacks such as FGSM may be more informative than stronger iterative ones when evaluated through a perturbation risk lens. Interestingly, PGN exhibits limited perturbation risk: despite outperforming PGD in the proposed analysis, the use of a Hessian regularization, known to enhance transferability across architectures (Ge et al., 2023), is effective mainly in high-$\kappa$ regimes. This questions the connection between stable loss curvature and robustness to statistically broad noisy perturbations.
Figure 5: Correlation between uniform project noise ($\mathcal{R}$ at $\kappa=0$) and $\mathcal{R}$ varying $\kappa$, using DN. (a) per-sample relation on ImageNet-Val samples for $\ell_{\infty}$ with $\varepsilon=16/255$. (b) Spearman coefficient $\rho_{s}$ vs. $\kappa$ for ImageNet-Val (top: $\varepsilon=16/255$ for $\ell_{\infty}$, $\varepsilon=2.0$ for $\ell_{2}$) and CIFAR-10 (bottom: $\varepsilon=8/255$ for $\ell_{\infty}$, $\varepsilon=1.0$ for $\ell_{2}$)
@@
Importantly, the reference concentration $\kappa^{*}=D^{1/4}$ (vertical dotted lines) reveals a substantial gap between $\bar{\mathcal{R}}$ and the ASR, even considering the DN attack. From a probabilistic perspective, this gap highlights that standard adversarial attacks primarily exploit brittle failure modes, which have limited relevance under realistic perturbation distributions (Fawzi et al., 2018; Carlini et al., 2019). Despite this common behavior, the DN attack exhibits a faster increase of $\bar{\mathcal{R}}$ with inital values of $\kappa$, indicating broader, noise-aligned failure regions at moderate concentrations. While not optimized for worst-case success, DN and FGSM remain more informative under perturbation risk evaluation, with similar, though more limited, trends observed for MIFGSM and PGN due to the use of noise during optimization.
@@ 6.3 Empirical validation of $\kappa^{*}$
The reference concentration $\kappa^{*}=D^{1/4}$ is introduced based on the analysis presented in Section 4 and Appendix B. To further validate this choice from an empirical perspective, we analyze the behavior of the directional perturbation risk for small values of $\kappa$, focusing on DN-based attacks, which, according to the previous benchmarks, achieve best performance in low $\kappa$ regimes. Specifically, we study the sample-wise correlation between $\mathcal{R}$ computed at fixed $\kappa$ values and that obtained under uniform projected noise ($\mathcal{R}$ with $\kappa=0$). As illustrated in Figure 5(a), for small concentration values (e.g., $\kappa=2$), the directional perturbation risk remains strongly correlated with the uniform-noise case; however, as $\kappa$ increases (e.g., $\kappa=10$), this similarity progressively decreases, indicating that the attack achieves high success rates often in highly directional regimes, while remaining largely ineffective under uniform noise.
To quantify this trend across multiple samples, Figure 5(b) reports the Spearman rank correlation coefficient $\rho_{s}$ between the $\mathcal{R}$ values at a given $\kappa$ (x-axis) and those at $\kappa=0$. Results are shown for both ImageNet-Val and CIFAR-10, and for $p=\infty$ and $p=2$, highlighting the effect of different input dimensionalities. In all settings, the rank correlation decreases sharply with $\kappa$, confirming that higher concentrations correspond to increasingly localized and less representative perturbation regimes. The reference value $\kappa^{*}=D^{1/4}$ ($\approx 19$ for ImageNet, $\approx 7$ for CIFAR-10) lies in a transitional region where rank correlation with uniform noise remains relatively high, yet begins to decay rapidly. This regime represents a meaningful trade-off: it preserves a strong connection to noise-aligned perturbations while avoiding high-$\kappa$ regime where the analysis becomes dominated by corner-case and too worst-case behavior.
@@ 6.4 On the reliability of model confidence
In addition to the analysis presented above, we investigate how model inference behavior relates to robustness across different $\kappa$ regimes, with the goal of identifying common characteristics of samples that are vulnerable to stochastic noise. Figure 7 reports the value of $\mathcal{R}$ for individual samples as a function of the model confidence on the corresponding clean input (x-axis), i.e., $\exp\big(f_{y}(x)\big)/{\sum_{c}\exp\big(f_{c}(x)\big)}.$ The analysis is conducted for different concentration values, $\kappa\in{5,20,50}$, using the DN attack, illustrating the shift from noise-like to highly directional regimes.
For both ResNet50 and ViT-B, a consistent pattern emerges. In the lowest concentration regime ($\kappa=5$), samples exhibiting high perturbation risk ($\mathcal{R}\approx 1$) are predominantly associated with low confidence scores. Conversely, samples predicted with high confidence remain largely robust, with $\mathcal{R}$ close to zero, despite being potentially vulnerable under classical worst-case adversarial evaluations. As $\kappa$ increases to moderate values (e.g., $\kappa=20\approx\kappa^{*}$), high perturbation risk gradually extends toward samples with higher confidence. This effect is more pronounced for ResNet models, while ViT architectures exhibit more stable reliability, with high-risk samples remaining largely concentrated among low-confidence predictions. Only at relatively large values of $\kappa$ (e.g., $\kappa\geq 50$, where perturbation risk becomes less informative) does the distribution of highly attacked samples ($\mathcal{R}$ high) significantly include high-confidence predictions.
The trend suggests that, under statistically plausible perturbation regimes, adversarial vulnerability could be strongly coupled with model uncertainty. This provides a more nuanced perspective on the relationship between confidence and robustness: while confidence is known to be unreliable under worst-case threat models (Grosse et al., 2018; Galil & El-Yaniv, 2021), it could remain informative for identifying samples that are realistically at risk under stochastic perturbations. This observation highlights a potential connection between robustness evaluation and uncertainty estimation, suggesting that uncertainty-aware analyses (Gawlikowski et al., 2023; Oberdiek et al., 2018) should not be dismissed based on their behavior under worst-case adversarial settings, and still play a meaningful role in safety-oriented model assessment. Results for additional models, exhibiting a similar trend, are reported in Appendix D.
@@ 7 Conclusion and Discussion
This work examines the role of adversarial attacks in robustness and safety evaluation. We introduce a probabilistic framework that bridges adversarial and noisy robustness through directional perturbation risk analysis, utilizing a concentration parameter to characterize the relationship between adversarial directions and noise-aligned perturbations. We then presented a directional noisy attack that explicitly targets statistically representative failure regions. Results show that many standard adversarial attacks achieve high success only in highly worst-case regimes, while simpler or noise-aware strategies can be more informative when robustness is evaluated from a perturbation risk perspective. These findings, rather than discouraging the use of attacks for this purpose, highlight important limitations of relying solely on attack success rates and invite practitioners to complement attack evaluation with perturbation risk analysis.
@@ Appendix A Sampling Directions on $\ell_{p}$-Spheres
Let $\xi\sim\mathcal{N}(0,I_{d})$ be a standard Gaussian vector in $\mathbb{R}^{d}$. We consider perturbations obtained by projection onto the $\ell_{p}$-sphere of radius $r$, $\eta=\Pi_{\|\cdot\|_{p}=r}(\xi).$
@@
Despite this lack of exact uniformity, Gaussian projection remains a practical and effective approximation in high dimensions. Indeed, by concentration of measure, $\|\xi\|_{p}^{p}=\sum_{i=1}^{d}|\xi_{i}|^{p}$ concentrates sharply around its expectation as $d$ increases. Consequently, normalization by $\|\xi\|_{p}$ primarily rescales $\xi$ by an almost deterministic factor, while preserving a diverse set of directions. As a result, $\xi/\|\xi\|_{p}$ provides a stable set of noise-aligned perturbations that is sufficient for estimating perturbation risk probabilities, while retaining exact uniformity in the $\ell_{2}$ case.
@@ Appendix B On the Scaling of $\kappa$ and $\kappa^{*}$
This places $\kappa^{\star}$ well within the noise-dominated regime, ensuring that samples remain broadly stochastic while introducing a controlled and measurable directional preference. In typical high-dimensional vision settings ($D\approx 10^{4}$$10^{6}$), this choice corresponds to moderate values of $\kappa$ that are empirically well below $\sqrt{D}$, yet sufficient to induce meaningful directional bias. Please note that $\kappa^{\star}=D^{1/4}$ is not intended to be an optimal or universal choice. Rather, it serves as a convenient reference point for the proposed analysis, lying well within the noise-dominated regime while still inducing a measurable directional bias. The conclusions of this work do not rely on the exact value of $\kappa^{\star}$, but instead on the qualitative behavior of the perturbation risk metric across low-$\kappa$ regimes. Similar trends are observed for a broad range of $\kappa$ values satisfying $\kappa\ll\sqrt{D}$, as further supported by the experimental analysis.
Table 2: Robustness scores under $\ell_{\infty}$ attacks ($\varepsilon\in\{4/255,8/255\}$) and $\ell_{2}$ attacks ($\varepsilon\in\{0.5,1.0\}$). Here, $\bar{\mathcal{R}}_{\kappa^{*}}$ denotes the average perturbation risk on the test set evaluated at $\kappa^{*}=D^{1/4}$, while $\kappa_{0.25}$ denotes the smallest concentration parameter $\kappa$ for which $\bar{\mathcal{R}}\geq 0.25$. ASR denotes the standard attack success rate, reported for comparison.
@@
Figure 9: Comparison of different adversarial attacks on the CIFAR10 tests set measured by the perturbation risk metric $\bar{\mathcal{R}}$ at small values of $\kappa$ for ResNet-32 and VGG-16 under $\ell_{2}$ perturbations. The dotted lines denote the reference point $\kappa^{*}=D^{1/4}$.